OpenStudy (anonymous):

A system of linear inequalities is shown below. x - y > 3 y + x ≤ 2 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

x - y > 3 put this in y = mx + b form -y > -x + 3 y < x - 3 slope = 1 y intercept is (0,-3) x intercept can be found by subbing in 0 for y x - y > 3 x - 0 > 3 x > 3 x intercept is (3,0) since the line does not have an equal sign, you will use a dashed line start at (0,-3) and since the slope is 1, go up one space, and to the right 1 space, then up 1 space, and to the right 1 space, keep doing this and you will cross the x axis at (3,0). and since the ine is less then, shading occurs below the line. y + x <= 2 y < = -x + 2 slope is 2 y intercept is (0,2) x intercept : y + x <= 2 0 + x <= 2 x <= 2 x intercept is (2,0) since there is an equal sign, the line will be solid start at (0,2) and since the slope is -1, go down 1, and to the right 1, then down 1, and to the right 1, and you will eventually cross the x axis at (2,0) and since the inequality sign is less then, shade below the line solution will be in the shaded regions that both lines share

OpenStudy (triciaal):

one method for each line set as an equation find y when x = 0 and x when y = 0 this will give 2 points connect to get the line with the eqn in the form y =mx + b replace with the inequality and shade that section the overlap with the shaded sections